You have 15 pennies in your pocket. Of those pennies, 3 are Canadian. Suppose you pick a penny out of your pocket at random. Find probability (not canadian). . A. 5 B. 1/5 C. 6/5 D. 4/5
The best and most correct answer among the choices provided by your question is the fourth choice or letter D. The probability is 4/5.
P ( canadian)= 3/15 = 1/5
On the other hand, P (not canadian) = 1 - 1/5 = 4/5
I hope my answer has come to your help. Thank you for posting your question here in Brainly.
Answer:
4/5
Step-by-step explanation:
just took the test on edge 2021